HELP!!! OK I tried to do this but it didn't work...
Problem: The length of a rectangle is 3 meters less than twice the width. The perimeter is 18 meters. Find the width.
So is it 18=2*2x-3*2x
THANK YOU SO MUCH AGAIN!

Answer by checkley71(8403)

L=2W-3
2(2W-3)+2W=18
4W-6+2W=18
6W=18+6
6W=24
W=24/6
W=4 FOR THE WIDTH.
L=2*4-3
L=8-3
L=5 FOR THE LENGTH.
PROOF
2*5+2*4=18
10+8=18
18=18